OpenStudy (anonymous):

2yx^0 is this whole thing become 1 then cancel out or just the x??????

jimthompson5910 (jim_thompson5910):

if the 0 is just an exponent for x, then x^0 = 1 the 2y doesn't change

jimthompson5910 (jim_thompson5910):

if you had (2yx)^0, then the whole thing would be 1

OpenStudy (anonymous):

alright thanks and btw it wasnt in parentheses so yea

jimthompson5910 (jim_thompson5910):

ok, then it's 2y

OpenStudy (anonymous):

alright what would happen next if it was (2yx^0 . x^3 y^2/2x^3 y^-2)-2

jimthompson5910 (jim_thompson5910):

The problem is this right? \[\Large \left(\frac{2yx^0*x^3y^2}{2x^3y^{-2}}\right)^{-2}\]

OpenStudy (anonymous):

yes lol

jimthompson5910 (jim_thompson5910):

what do you get when you simplify the numerator?

OpenStudy (anonymous):

2y .x^3 y^2

jimthompson5910 (jim_thompson5910):

what is y times y^2 equal to?

OpenStudy (anonymous):

y^3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

we now have \[\Large \left(\frac{2x^3y^3}{2x^3y^{-2}}\right)^{-2}\]

jimthompson5910 (jim_thompson5910):

now try to simplify the fraction inside the parenthesis (ignore the negative -2 exponent on the very outside)

OpenStudy (anonymous):

would it be 1y^5?

OpenStudy (anonymous):

or just y^5

jimthompson5910 (jim_thompson5910):

good, \[\Large \left(\frac{2x^3y^3}{2x^3y^{-2}}\right)^{-2} = (y^5)^{-2}\]

jimthompson5910 (jim_thompson5910):

when you raise an exponential expression to another exponent, you multiply the exponents

OpenStudy (anonymous):

so -25? but wouldnt u flip the equation because its a negative exponent?

jimthompson5910 (jim_thompson5910):

5 times -2 is not -25

OpenStudy (anonymous):

(1/y^5)^2

jimthompson5910 (jim_thompson5910):

yes you did a valid move, so, \[\Large \left(\frac{1}{y^5}\right)^{2} = \frac{1^2}{(y^5)^2} = \frac{1}{y^{10}}\]

jimthompson5910 (jim_thompson5910):

be careful and notice how I got 10 as the exponent and NOT 25

jimthompson5910 (jim_thompson5910):

you're tempted to say 5^2 = 25 but it should be 5*2 = 10

OpenStudy (anonymous):

yea lol

OpenStudy (anonymous):

4^b-3=64 idk how to do this apparently i have to make it something like 4^b-3=4^4?

jimthompson5910 (jim_thompson5910):

b-3 is the exponent?

OpenStudy (anonymous):

yes